# Hyper-Hermitian Weyl Double Copy

### v1

@@ 3 Hyper-Hermitian and Hyper-Kählerian Weyl Double Copy
In the present section we study a certain class of hyper-Hermitian metrics called elementary states. After some preliminary material on the Weyl double copy, we introduce the definitions of the elementary states. We apply the Weyl double copy to perform a systematic analysis of the double ans single copies of these elementary states.
@@ 3.1 Overview of the Weyl double copy
In the present subsection, we provide the procedure for constructing double copy examples. In this and subsequent sections, for notation and conventions, we follow the Penrose and Rindler classical book [61]. Thus we first recall some general aspects of the Weyl double copy following mainly [23]. Consider the Weyl curvature $\Psi_{ABCD}$ with the symmetry property $\Psi_{ABCD}=\Psi_{(ABCD)}$, where $A=(0,1)$ is a spinorial index. In the Weyl double copy, the Weyl curvature can be written as
@@
In general the Maxwell field strength spinors $\Phi^{(1)}_{AB}$ and $\Phi^{(2)}_{AB}$ are different, in the single copy they are equal. Then eq. (17) expresses the relationship between Einstein and Maxwell solutions. We propose a family of $\Phi$ functions that we will use in the rest of the paper which is given by
@@
for all $r\in\mathbb{R}$ and where $A$, $B_{r}$ and $C_{r}$ are constants. Moreover, $\Phi$ is a solution of the wave equation
@@
We note that we can change $u\leftrightarrow v$, $x\leftrightarrow y$ in $\Phi$ and it will still satisfy eq. (19). These types of properties are observed for example in [49]. To obtain the Maxwell tensor field we use the spinorial formalism. We have that the ’spinorial vierbein’ $\sigma^{\mu}_{AA^{\prime}}$ satisfies the relation
@@
Now we are going to use this object to write down any tensor with Greek indices, $V_{AA^{\prime}}=\sigma^{\mu}_{AA^{\prime}}V_{\mu}$, this also applies to replace Greek indices by Latin indices. We have
@@
Let $\xi^{A}=(1,Z)$, then $\Psi_{ABCD}\xi^{A}\xi^{B}\xi^{C}\xi^{D}$ is a degree 4 polynomial in variable $Z$. Therefore it can be written as
@@
where $P_{i}(Z)$ is a monomial and $\alpha(x^{a})$ is a function of the coordinates $x^{a}=\{x,y,u,v\}$. On the other hand, $\Phi_{AB}\xi^{A}\xi^{B}$ is also a polynomial of $Z$, but up to degree 2, thus we have
@@
where $i\in\{1,2\}$, and $h_{1}(x^{a})$, $h_{2}(x^{a})$ are also functions of the coordinates. Using (25) we get
@@
then, from (24) and (26), we have the following implication
@@
On the other hand, we have that any Maxwell spinor can be written as the symmetrized product of two spinors $\textbf{r}_{A}$ and $\textbf{s}_{A}$
@@
Then in terms of $Z$ from $\xi^{A}=(1,Z)$ we have
@@
Comparing equations (25) and (30), we found the coefficients of eq. (28).
Now in spinorial formalism, source-free Maxwell’s equations are given by [61]
@@
Equations (32) and (33) are differential equations for $h_{i}=h_{i}(x,y,u,v)$ and they are equal if $\Phi^{A}{}_{B}$’s are equal.
@@
we use equation (27) in order to find $h_{i}(x^{a})$ and it is possible to verify that they satisfy the equation (31), otherwise we solve differential equation (34) to find $h_{i}(x^{a})$.
@@ 3.2 Hyper-Hermitian Weyl double copy: Preliminaries
In this subsection, we give a brief overview of the family of solutions called elementary states introduced in ref. [57]. Let $x^{AA^{\prime}}=(x^{A},u^{A})$ be local null coordinates on a complex four-manifold $\mathcal{M}$ and let $\Theta_{A}$ be a pair of complex valued functions (potentials) on $\cal{M}$, we take $x^{A}=(y,-x)$ and $u^{A}=(u,v)$. We write down the equations (16) once again for convenience
@@
In the present article we will be interested in a family of solutions to eq, (16) of the hyper-Hermitian metrics called elementary states [57], given by
@@
Then the class of solutions of our interest in the present article for the components of the $F_{A}$ functions will be of the polynomial form
@@
where $a_{i}$ and $b_{i}$ with $i\in\{1,2,3,4\}$ are constants. In order for eqs. (36) with (38) satisfy $\partial^{2}\Theta_{C}=0$, the parameters of that solution must take certain values. For the equation $\partial^{2}\Theta_{C}=0$ to be satisfied, is necessary to choose at least one of the three conditions from the set of parameters $\{a_{1}=0,a_{4}=a_{2}+a_{3},a_{4}=-1\}$ and at least one condition from the set $\{b_{1}=0,b_{4}=b_{2}+b_{3},b_{4}=-1\}$. Thus in total, we will have nine cases. For instance, the selection of $a_{1}=0$ and $b_{1}=0$ yields the trivial solution, the other eight cases give nontrivial ones. However, three cases will be repeated due to the symmetry under the interchange of parameters $a\leftrightarrow b$. One of these cases is $\{a_{1}=0,b_{4}=-1\}$ and $\{a_{4}=-1,b_{1}=0\}$.
@@
With (38) the contracted Weyl curvature spinor is given by
@@
where there is no restriction on the set of parameters. Thus for generic hyper-Hermitian manifolds, we can see that the Weyl curvature has tree roots equal and the other different and consequently we get a type III metric of the Petrov classification.
@@
On the other hand, starting also from the general case with unrestricted parameters we establish the conditions under which Weyl’s spinor leads to type N solutions. Thus one has
@@
Recall that a solution of vacuum Einstein equations is type N if it has four principal null directions equal, and from eqs. (43) and (44), it happens if the following relations are fulfilled
@@
Those conditions are precisely the same that we found in eq. (42). Thus, we then have that the result elementary states have to be type N is equivalent to the condition elementary states have to be hyper-Kähler . We summarize these results in the table 1 in which we omit the parameter values that give us the same metric, for example, $a_{1}=0$ and $b_{4}=b_{2}+b_{3}$, give the same metric as $a_{4}=-1$ and $b_{4}=b_{2}+b_{3}$.
@@
Table 1: In the table it is depicted the parameters that allow to build elementary states and how they are ordered according to how they give rise to hyper-Hermitian or hyper-Kähler solutions. For the hyper-Hermitian case they generically lead to type III solutions and the hyper-Kähler give type N solutions. For the hyper-Kähler case we restricted the allowed of the parameters from eqs. (36) and (38), and then impose separately the implementation of conditions $\partial^{2}\Theta_{A}=0$ and $\partial_{A}\Theta^{A}$=0.
@@ 3.3 Hyper-Hermitian Weyl double copy: Generic case
In the present subsection, we provide a series of examples of hyper-Kählerian and hyper-Hermitian double copy correspondence. We start from the mentioned case of elementary states encoded in the family of solutions given by eqs. (36) and (38)
@@
The contracted spinorial Weyl curvature can be written by using (23) as follows
@@
Moreover, we know that $\Phi^{(1/2)}_{AB}\xi^{A}\xi^{B}$ is a polynomial in the variable $Z$ up to degree 2, but in principle, we do not know how to find the factor $\alpha(x^{a})$ between the two contracted Maxwell spinors. Therefore the field strength spinors are
@@
where $h_{i}(x^{a})$ represent our lack of knowledge about the factor $\alpha(x^{a})$. Now from the comparison between equations (28) and (30) we can find the spinor entries
@@
Since the polynomial $(u-vZ)^{2}$ is of the same form as for the Maxwell spinor in Sparling-Tod double copy (81) (see below), we establish that
@@
for all $r\in\mathbb{R}$ and where $A$, $B_{r}$ and $C_{r}$ are constants as we mentioned previously. Finally, the explicit form of Maxwell’s spinor is written as
@@
Equation (56) satisfies (31). Analogously the field strength is given by
@@
One can check that eq. (57) satisfies (34), where the current $J^{AA^{\prime}}$ is not conserved in the following form
@@
Thus generically in the hyper-Hermitian case, besides the existence of two different Maxwell’s spinors, one of them will be associated with a source current and the other will be sourceless. This procedure will be the same for all the following examples that will be described in the next section. This is, therefore, one of our main results: for the Weyl double copy in Hermitian manifolds and elementary states in general, it holds that
@@
where $\Psi_{ABCD}$ is given by eq. (47), $\Phi$ correspond to (55) and $\Phi_{AB}^{(1)}$ and $\Phi_{AB}^{(2)}$ are given by (56), (57) respectively with $\{a_{1}=0$, $b_{4}=b_{2}+b_{3}\}$ given one metric or $\{a_{4}=a_{2}+a_{3}$, $b_{4}=b_{2}+b_{3}\}$ given other metric. We called these results generic hyper-hermitian double copy. We note that, since the equation $\partial^{2}\Theta_{A}=0$ is linear, we can take linear combinations of the potential (46) without any issue, which highlights the generality of the result.
@@ 4 Examples
In this Section we will give several examples regarding our results of the previous section. We will discuss the explicit Weyl double copies for five examples. Two of them will lead to the hyper-Kähler double copy case and the rest will be genuine hyper-Hermitian.
@@ 4.1 Hyper-Kähler double copy: Generic case
In what follows we will consider the solutions $\Theta_{A}$ of the form given by eq. (46) with $a_{4}=a_{2}+a_{3}$, $b_{4}=b_{2}+b_{3}$, $b_{2}=a_{2}+1$, $a_{3}=b_{3}+1$, $a_{1}=-b_{1}$. Thus we start from the following pair of potentials
@@
The metric is type N, then the field strength’s spinor has the form
@@
thus in this case both Maxwell’s spinors are equal. Now imposing eq. (27) the function $h(x^{a})$ can be written in the form
@@
where $\Phi$ is given by (18) and $c_{1}\in\mathbb{R}$ is absorbed in the constants of $\Phi$. Writing explicitly the only Maxwell spinor we have
@@
Figure 1: In this figure the electric field obtained from Maxwell field strength (71) of subsection 4.1. Some parameters are selected just with $c_{1}=A=r=b_{2}=1$ and $t=B_{r}=C_{r}=b_{3}=z=0$.
To visualize the solution we plot the electric field obtained from eq. (71) in figure 1 in terms of the Minkowskian coordinates $(x^{0},x^{1},x^{2},x^{3})$ which are defined by
@@ 4.2 Sparling-Tod Double Copy
The Sparling-Tod metric can be obtained from (64) for the following specific values of parameters: $b_{3}=0$, $b_{2}=1$, $a_{1}=-1$. Thus, the metric is expressed as
@@
In this case knowing that this solution is of type N, the contracted Weyl curvature reads
@@
From eq. (24) one can be read off the polynomials
@@
Then we find Maxwell’s spinors in the following form
@@
Explicitly we find that there is only one Maxwell’s spinor and it is given by
@@
where, with the aid of equation (27), we obtain
@@
Here $\Phi$ is given by equation (18) and it reproduces Sparling-Tod metric when $\Phi=(ux+yv)^{-1}$. With these ingredients is possible to find the explicit form of Maxwell’s spinor as follows:
@@
for all $r\in\mathbb{R}$ and where $A$, $B_{r}$ and $C_{r}$ are constants as we already mentioned. These results with different families of solutions and only one metric were obtained similarly in ref. [20]. We plot in Figure 2 the equation (84) with different values of the parameters.
(a) Here the selected parameters are: $A=1$, $B_{r}=0$, $C_{r}=0$, $z=0$.
(b) The used parameters are: $A=1$, $B_{r}=1$, $C_{r}=1$, $r=2$, $z=1$ and $t=0$.
Figure 2: In these figures we plot Maxwell’s spinor (84) corresponding to Sparling-Tod’s electric field with different values of parameters in $\Phi$.
So far, the results we have obtained have given us metrics of type N, this is expected since the two metrics are hyper-Kähler. In the following examples, we see metrics that are not hyper-Kähler and with leading null addresses that have a multiplicity different from 4, all these results are summarized in the tables 2 and 3.

### v2

@@ 3 Hyper-Hermitian and Hyper-Kählerian Weyl Double Copy
In the present section, we study a certain class of hyper-Hermitian metrics called elementary states. After some preliminary material on the Weyl double copy, we introduce useful definitions. We apply the Weyl double copy to perform a systematic analysis of the double and single copies of these elementary states.
@@ 3.1 Overview of the Weyl double copy
In the present subsection, we provide the procedure for constructing double-copy examples. In this and subsequent sections, for notation and conventions, we follow the Penrose and Rindler classic book [64]. Thus, we first recall some general aspects of the Weyl double copy following mainly ref. [25]. Consider the Weyl curvature $\Psi_{ABCD}$ with the symmetry property $\Psi_{ABCD}=\Psi_{(ABCD)}$, where $A=(0,1)$ is a spinorial index. In the Weyl double copy, the Weyl curvature can be written as
@@
In general, the Maxwell field strength spinors $\Phi^{(1)}{AB}$ and $\Phi^{(2)}{AB}$ are different; in the single copy, there is only one Maxwell spinor. Then, eq. (17) expresses the relationship between Einstein and Maxwell solutions. We propose a family of $\Phi$ functions that we will use in the rest of the paper which is given by
@@
for all $r\in\mathbb{R}$, where $A$, $B_{r}$ and $C_{r}$ are constants. Moreover, $\Phi$ is a solution of the wave equation
@@
We note that we can change $u\leftrightarrow v$, $x\leftrightarrow y$ in $\Phi$ and it will still satisfy eq. (19). These types of properties are observed for example in ref. [52]. To obtain the Maxwell tensor field we use the spinorial formalism. We have that the ’spinorial vierbein’ $\sigma^{\mu}_{AA^{\prime}}$ satisfies the relation
@@
Now, we are going to use this object to write down any tensor with Greek indices, $V_{AA^{\prime}}=\sigma^{\mu}_{AA^{\prime}}V_{\mu}$. This also applies to replacing Greek indices with Latin indices. We have
@@
Let $\xi^{A}=(1,Z)$, then $\Psi_{ABCD}\xi^{A}\xi^{B}\xi^{C}\xi^{D}$ is a degree 4 polynomial in variable $Z$. Therefore, it can be written as
@@
where $P_{i}(Z)$ is a monomial and $\alpha(x^{a})$ is a function of the coordinates $x^{a}=\{x,y,u,v\}$. On the other hand, $\Phi_{AB}\xi^{A}\xi^{B}$ is also a polynomial in $Z$, but up to degree 2. Thus, we have
@@
where $i\in\{1,2\}$, and $h_{1}(x^{a})$, $h_{2}(x^{a})$ are also functions of the coordinates. Using eq. (25) we get
@@
then, from eqs. (24) and (26), we have the following implication
@@
On the other hand, any Maxwell spinor can be written as the symmetrized product of two spinors $\textbf{r}_{A}$ and $\textbf{s}_{A}$
@@
Then, in terms of $Z$ from $\xi^{A}=(1,Z)$, we have
@@
Comparing eqs. (25) and (30), we find the coefficients of eq. (28).
Now, in spinorial formalism, source-free Maxwell’s equations are given by [64]
@@
Eqs. (32) and (33) are differential equations for $h_{i}=h_{i}(x,y,u,v)$ and they are equal if $\Phi^{A}{}_{B}$’s are equal.
@@
we use eq. (27) in order to find $h_{i}(x^{a})$ and it is possible to verify that they satisfy the eq. (31), otherwise, we solve the differential eq. (34) to find $h_{i}(x^{a})$.
@@ 3.2 Hyper-Hermitian Weyl double copy: Preliminaries
In this subsection, we give a brief overview of the family of solutions called elementary states introduced in ref. [60]. Let $x^{AA^{\prime}}=(x^{A},u^{A})$ be local null coordinates on a complex four-manifold $\mathcal{M}$ and let $\Theta_{A}$ be a pair of complex-valued functions (potentials) on $\cal{M}$, we take $x^{A}=(y,-x)$ and $u^{A}=(u,v)$. We write down the eqs. (16) once again for convenience
@@
In the present article, we will be interested in a family of solutions to eq. (16) of the hyper-Hermitian metrics called elementary states [60], given by
@@
Then, the class of solutions of our interest in the present article for the components of the $F_{A}$ functions will be of the polynomial form
@@
where $a_{i}$ and $b_{i}$, with $i\in\{1,2,3,4\}$, are constants. In order for eqs. (36) with (38) satisfy $\partial^{2}\Theta_{C}=0$, the parameters of that solution must take certain values. For the equation $\partial^{2}\Theta_{C}=0$ to be satisfied, it is necessary to choose at least one of the three conditions from the set of parameters $\{a_{1}=0,a_{4}=a_{2}+a_{3},a_{4}=-1\}$ and at least one condition from the set $\{b_{1}=0,b_{4}=b_{2}+b_{3},b_{4}=-1\}$. Thus, in total, we will have nine cases. For instance, the selection of $a_{1}=0$ and $b_{1}=0$ yields the trivial solution, the other eight cases give nontrivial ones. However, three cases will be repeated due to the symmetry under the interchange of parameters $a\leftrightarrow b$. One of these cases is $\{a_{1}=0,b_{4}=-1\}$ and $\{a_{4}=-1,b_{1}=0\}$.
@@
With eq. (38) the contracted Weyl curvature spinor is given by
@@
where there is no restriction on the set of parameters. Thus, for generic hyper-Hermitian manifolds, we can see that the Weyl curvature has three equal roots and the other different and consequently we get a type III metric of the Petrov classification.
@@
On the other hand, starting also from the general case with unrestricted parameters we establish the conditions under which Weyl spinor leads to type N solutions. Thus one has
@@
Recall that a solution of the vacuum Einstein equations is type N metric if it has four principal null directions equal, and from eqs. (43) and (44). It happens if the following relations are fulfilled
@@
Those conditions are precisely the same that we found in eq. (42). Thus, we have that the result elementary states have to Petrov type N metric is equivalent to the condition elementary states have to be hyper-Kähler. We summarize these results in Table 1 in which we omit the parameter values that give us the same metric, for example, $a_{1}=0$ and $b_{4}=b_{2}+b_{3}$, give the same metric as $a_{4}=-1$ and $b_{4}=b_{2}+b_{3}$.
@@
Table 1: In the table is depicted the parameters that allow to build elementary states and how they are ordered according to how they give rise to hyper-Hermitian or hyper-Kähler solutions. For the hyper-Hermitian case they generically lead to type III solutions and the hyper-Kähler give type N solutions. For the hyper-Kähler case we restricted the allowed of the parameters from eqs. (36) and (38), and then impose separately the implementation of conditions $\partial^{2}\Theta_{A}=0$ and $\partial_{A}\Theta^{A}$=0.
@@ 3.3 Hyper-Hermitian Weyl double copy: Generic case
In the present subsection, we provide a series of examples of the hyper-Kählerian and the hyper-Hermitian double copy correspondence. We start from the mentioned case of elementary states encoded in the family of solutions given by eqs. (36) and (38)
@@
The contracted spinorial Weyl curvature can be written by using eq. (23) as follows
@@
Moreover, we know that $\Phi^{(1/2)}_{AB}\xi^{A}\xi^{B}$ is a polynomial in the variable $Z$ up to degree 2, but in principle, we do not know how to find the factor $\alpha(x^{a})$ between the two contracted Maxwell spinors. Therefore, the field strength spinors are
@@
where $h_{i}(x^{a})$ represent our lack of knowledge about the factor $\alpha(x^{a})$. Now, from the comparison between eqs. (28) and (30), we can find the spinor entries
@@
Since the polynomial $(u-vZ)^{2}$ is of the same form as for the Maxwell spinor in Sparling-Tod double copy in eq. (81) (see below), we establish that
@@
for all $r\in\mathbb{R}$ and where $A$, $B_{r}$ and $C_{r}$ are constants as we mentioned previously. Finally, the explicit form of the Maxwell spinor is written as
@@
Eq. (56) satisfies eq. (31). Analogously, the field strength spinor $\Phi_{AB}^{(2)}$ is given by
@@
One can check that eq. (57) satisfies eq. (34), where the current $J^{AA^{\prime}}$ is not conserved in the following form
@@
Thus, generically in the hyper-Hermitian case, besides the existence of two different Maxwell spinors, one of them will be associated with a source current and the other will be sourceless. The process in which the Weyl double copy includes a source in the gauge sector has been studied in refs. [19, 20]. We reproduce its result, where one Maxwell spinor is sourceless, and at least one Maxwell spinor has a source. This procedure will be the same for all the following examples that will be described in the next section. This is, therefore, one of our main results: for the Weyl double copy in Hermitian manifolds and elementary states in general, it holds that
@@
where $\Psi_{ABCD}$ is given by eq. (47), $\Phi$ corresponds to eq. (55) and $\Phi_{AB}^{(1)}$ and $\Phi_{AB}^{(2)}$ are given by eqs. (56), (57) respectively with $\{a_{1}=0$, $b_{4}=b_{2}+b_{3}\}$ given one metric or $\{a_{4}=a_{2}+a_{3}$, $b_{4}=b_{2}+b_{3}\}$ given other metric. We called these results generic hyper-hermitian double copy. We note that, since the equation $\partial^{2}\Theta_{A}=0$ is linear, we can take linear combinations of the potential in eq. (46) without any issue, which highlights the generality of the result.
@@ 4 Examples
In this Section, we will give several examples regarding our results of the previous section. We will discuss the explicit Weyl double copies for five examples. Two of them will lead to the hyper-Kähler double copy case and the rest will be genuine hyper-Hermitian.
@@ 4.1 Hyper-Kähler double copy: Generic case
In what follows, we will consider the solutions $\Theta_{A}$ of the form given by eq. (46) with $a_{4}=a_{2}+a_{3}$, $b_{4}=b_{2}+b_{3}$, $b_{2}=a_{2}+1$, $a_{3}=b_{3}+1$, $a_{1}=-b_{1}$. Thus we start from the following pair of potentials
@@
We have a type N metric, then the field strength spinor has the form
@@
thus, in this case, both Maxwell spinors are equal. Now, imposing eq. (27) the function $h(x^{a})$ can be written in the form
@@
where $\Phi$ is given by eq. (18) and $c_{1}\in\mathbb{R}$ is absorbed in the constants of $\Phi$. Writing explicitly the only Maxwell spinor we have
@@
Figure 1: In this figure the electric field obtained from the Maxwell field strength in eq. (71) of subsection 4.1. Some parameters are selected just with $c_{1}=A=r=b_{2}=1$ and $t=B_{r}=C_{r}=b_{3}=z=0$.
To visualize the solution we plot the electric field obtained from eq. (71) in Figure 1, in terms of Minkowskian coordinates $(x^{0},x^{1},x^{2},x^{3})$ which are defined by
@@ 4.2 Sparling-Tod Double Copy
The Sparling-Tod metric can be obtained from eq. (64) for the following specific values of parameters: $b_{3}=0$, $b_{2}=1$, $a_{1}=-1$. Thus, the metric is expressed as
@@
In this case knowing that this solution is type N metric, the contracted Weyl curvature reads
@@
From eq. (24), one can be read off the polynomials
@@
Then, we find the Maxwell spinors in the following form
@@
Explicitly, we find that there is only one Maxwell spinor, and it is given by
@@
where, with the aid of eq. (27), we obtain
@@
Here $\Phi$ is given by eq. (18) and it reproduces Sparling-Tod metric when $\Phi=(ux+yv)^{-1}$. With these ingredients, it is possible to find the explicit form of Maxwell spinor as follows:
@@
for all $r\in\mathbb{R}$ and where $A$, $B_{r}$ and $C_{r}$ are constants as we already mentioned. These results with different families of solutions and only one metric were obtained similarly in ref. [22]. An example of a type D metric is the Eguchi-Hanson metric, as described in ref. [60], which is obtained with an Euclidean slice. We plot in Figure 2 eq. (84) with different values of the parameters.
(a) Here, the selected parameters are: $A=1$, $B_{r}=0$, $C_{r}=0$, $z=0$.
(b) The parameters used are: $A=1$, $B_{r}=1$, $C_{r}=1$, $r=2$, $z=1$ and $t=0$.
Figure 2: In these figures, we plot the Maxwell spinor in eq. (84) corresponding to Sparling-Tod’s electric field with different values of parameters in $\Phi$.
So far, the results we have obtained are type N metrics. This is expected since the two metrics are hyper-Kähler. In the following examples, we see metrics that are not hyper-Kähler and with leading null directions that have a multiplicity different from 4, all these results are summarized in the Tables 2 and 3.